OpenStudy (anonymous):

find the intercepts for the graph of the equation given 3x-5y=30

OpenStudy (anonymous):

whats the x and y intercept

OpenStudy (anonymous):

3x-5y-30=0 x-intercept (y=0) 3x-5(0)-30=0 3x=30 x=10 y-Intercept(x=0) 3(0)-5y-30=0 5y=-30 y=-6

OpenStudy (anonymous):

awwwwww thanx ;)

OpenStudy (anonymous):

np

OpenStudy (anonymous):

gave u a medal:)